OpenStudy (anonymous):

Use Lagrange multipliers to find the exact maximum and minimum values of f(x,y,z) = x^2+y^2+z^2 subject to the constraints x+y+2z=2 and x^2+y^2-z=0

OpenStudy (anonymous):

2x=λ + 2xμ, 2y = 2λ + 2yμ, 2z =2λ - μ, I can't figure out how to solve this system of equations.